Typically a bullet leaves a standard 45 caliber pistol at 262 m/s. If it takes 1 millisecond to transverse the barrel, determine the average acceleration experienced by the 16.2-gram bullet within the gun's barrel and then compute the average force exerted on it.
1
Expert's answer
2020-11-12T09:35:05-0500

By definition, the acceleration is


"a = \dfrac{\Delta v}{t}"

where "\Delta v = 262m/s-0m/s = 262m/s" is the change in speed and "t = 1ms = 10^{-6}s" is the time. Thus


"a = \dfrac{\Delta v}{t} = \dfrac{262m/s}{10^{-6}s} = 2.62\times 10^8m/s^2"

According to the Newton's second law, the force is


"F = ma = 16.2\times 10^{-3}kg\cdot 2.62\times 10^8m/s^2 = 4.2444\times 10^6N"

Answer. "a = 2.62\times 10^8m/s^2" and "F = 4.2444\times 10^6N".